OpenStudy (anonymous):

Can someone answer question 8 october november 2013 9701 paper 1/12 AS level chemistry

OpenStudy (abmon98):

Note % as mass A) 89.8*2/100 + 10.2*4/100=2.192 grams B) 80*2/100 + 19*4/100 + 16*1/100=2.52 grams C) 96.3*2/100 + 3.25*4/100 + 0.45*16/100=2.128 grams D) 82.5*2/100 + 15.2*4/100 + 2.3*16/100=2.626 grams Density=Mass/Volume 2.626/1=2.626g/dm^3